A body moving with simple harmonic motion has an amplitude of 10cm and a frequency of 100HZ, find (a) the period of oscillation (b) the acceleration at the maximum displacement (c) the velocity at the centre of motion.
1
Expert's answer
2021-10-12T16:23:02-0400

"s=A\sin(\omega t)=0.1\sin(2\pi\nu t)"


(a) "T=1/\nu=1/100=0.01\ (s)"


(b) "|a|=0.1\cdot(2\pi\nu)^2=0.1\cdot(2\cdot3.14\cdot100)^2=39438.4\ (m/s^2)"


(c) "|v|=0.1\cdot(2\pi\nu)=0.1\cdot(2\cdot3.14\cdot100)=6.28\ (m/s)"